A school has a ratio of men to women of 4:9 if there are 728 students at the school, how many are men?
Answer by ikleyn(52781)

.

As the condition says, there are 4+9 = 13 equal parts (of students), so each part has 728/13 = 56 students.


Therefore, the number of women is 4*56 = 224.     ANSWER